OpenStudy (anonymous):

solve for x: 855=5x^3

OpenStudy (anonymous):

whr r u stuck

OpenStudy (anonymous):

on finding x...

OpenStudy (anonymous):

We begin: \[ 855=5x^3\\ 171=x^3\\ \sqrt[3]{171}=\sqrt[3]{x^3}=x \]

OpenStudy (anonymous):

which will give u?

OpenStudy (anonymous):

the cube root of 171